# Supersymmetric BMS4 Algebras Revisited: Electric/Magnetic Superalgebras and Free Field Realization

### v1

@@ 2.2 Electric super BMS4 algebra
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},1)}_{r,\bar{r}},Q^{(1,\frac{1}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},1)}_{r,\bar{r}},Q^{(1,\frac{1}{2})}_{s,\bar{s}}\}=\left(r-\frac{1}{2}s\right)\left(\frac{1}{2}\bar{r}-\bar{s}\right)M_{r+s,\bar{r}+\bar{s}},$ (2.9w)
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(1,\frac{3}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(1,\frac{3}{2})}_{s,\bar{s}}\}=\left(r-\frac{1}{2}s\right)\frac{3}{2}\bar{r}M_{r+s,\bar{r}+\bar{s}},$ (2.9x)
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{3}{2},0)}_{r,\bar{r}},Q^{(0,\frac{3}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{3}{2},0)}_{r,\bar{r}},Q^{(0,\frac{3}{2})}_{s,\bar{s}}\}=\frac{9}{4}r\bar{s}M_{r+s,\bar{r}+\bar{s}}.$ (2.9y)
holomorphic (2.9b), anti-holomorphic (LABEL:eq:TensorProductOfVirReprsSolution3):
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(1,\frac{5}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(1,\frac{5}{2})}_{s,\bar{s}}\}=\left(r-\frac{1}{2}s\right)\bar{r}\left(\frac{5}{2}\bar{r}+\bar{s}\right)M_{r+s,\bar{r}+\bar{s}},$ (2.9z)
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{3}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{3}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{s,\bar{s}}\}=\frac{3}{2}s\bar{r}\left(\frac{5}{2}\bar{r}+\bar{s}\right)M_{r+s,\bar{r}+\bar{s}}.$ (2.9aa)
holomorphic (LABEL:eq:TensorProductOfVirReprsSolution3), anti-holomorphic (LABEL:eq:TensorProductOfVirReprsSolution3):
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{5}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{5}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{s,\bar{s}}\}=s\left(r+\frac{5}{2}s\right)\bar{r}\left(\frac{5}{2}\bar{r}+\bar{s}\right)M_{r+s,\bar{r}+\bar{s}}.$ (2.9ab)
All of the above algebras contain finite-dimensional subalgebras. However, except for the case (2.9s), the corresponding finite-dimensional supercharges necessarily include spin-$\frac{3}{2}$ or spin-$\frac{5}{2}$ components. Such higher-spin fermionic generators are not compatible with the Haag–Łopuszański–Sohnius theorem, which allows only spin-$\frac{1}{2}$ supercharges in a consistent extension of the Poincaré algebra. Therefore, the only physically admissible super Poincaré algebra is the finite-dimensional subalgebra of (2.9s).
@@ 2.3 Magnetic chiral super BMS4 algebras
$M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=\left(\frac{1}{2}r-\frac{1}{2}s\right)Q^{(0,\frac{1}{2})}_{r+s,\bar{r}},\quad\{Q^{(\frac{1}{2},0)}_{r},Q^{(\frac{1}{2},0)}_{s}\}=L_{r+s},\quad\{Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{s,\bar{s}}\}=M_{r+s,\bar{s}},$ (2.9ac)
@@
$M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=Q^{(1,\frac{1}{2})}_{r+s,\bar{r}},\quad\{Q^{(\frac{1}{2},0)}_{r},Q^{(\frac{1}{2},0)}_{s}\}=L_{r+s},\quad\{Q^{(\frac{1}{2},0)}_{r},Q^{(1,\frac{1}{2})}_{s,\bar{s}}\}=-\left(r-\frac{1}{2}s\right)M_{r+s,\bar{s}},$ (2.9ad)
In Type I, $M,Q^{(\frac{1}{2},0)}]\propto Q^{(0,\frac{1}{2})}$ satisfies (2.9b), whereas in Type II, $M,Q^{(\frac{1}{2},0)}]\propto Q^{(1,\frac{1}{2})}$ satisfies (2.9a). Unlike the electric case, the supercharges $Q^{(\frac{1}{2},0)}_{r}$ in (2.9ac) and (2.9ad) are not extended to $Q^{(\frac{1}{2},0)}_{r,\bar{r}}$, thereby avoiding the introduction of nonphysical generators $L_{m,\bar{m}}$. In contrast, the holomorphic index $r$ of $Q^{(0,\frac{1}{2})}_{r,\bar{r}}$ in (2.9ac) is required for the nonvanishing commutator $M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=Q^{(1,\frac{1}{2})}_{r+s,\bar{r}}$.
@@
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},\tilde{M}_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{r,\bar{r}},Q^{(1,\frac{1}{2})}_{r,\bar{r}}\}$, (2.9ae)
@@
Observing (2.9ac), we note that $Q^{(\frac{1}{2},0)}$ and $Q^{(0,\frac{1}{2})}$ are not related by conjugation. This makes it clear that both (2.9ac) and (2.9ad) describe chiral superalgebras, as only the holomorphic Virasoro algebra $L_{n}$ appears in anticommutators of super charges. For this reason, (2.9ac) and (2.9ad) are considered as holomorphic in this paper. Anti-holomorphic counterparts $\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(\frac{1}{2},0)}_{r,\bar{r}}\}$ and $\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}}\}$ can be defined through conjugation, yielding super BMS4 algebra structures analogous to (2.9ac) and (2.9ad).
For a super BMS4 algebra containing both holomorphic and anti-holomorphic sectors, simply including the corresponding supercharges is insufficient. For instance, consider holomorphic supercharges $\{Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{r,\bar{r}}\}$ and anti-holomorphic supercharges $\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(\frac{1}{2},0)}_{r,\bar{r}}\}$, both of Type I as in (2.9ac). The Jacobi identity for $\{Q^{(\frac{1}{2},0)},Q^{(\frac{1}{2},0)},\bar{Q}^{(\frac{1}{2},0)}\}$ is satisfied only if a bosonic generator $R^{(0,0)}$ is introduced, such that
$\{Q^{(\frac{1}{2},0)},\bar{Q}^{(\frac{1}{2},0)}\}\propto R^{(0,0)}.$ (2.9af)
@@
$[M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=\left(\frac{1}{2}r-\frac{1}{2}s\right)Q^{(0,\frac{1}{2})}_{r+s,\bar{r}},\quad[M_{r,\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{s}}]=\left(\frac{1}{2}\bar{r}-\frac{1}{2}\bar{s}\right)\bar{Q}^{(\frac{1}{2},0)}_{r,\bar{r}+\bar{s}}$, (2.9ag)
@@
$[M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=Q^{(1,\frac{1}{2})}_{r+s,\bar{r}},\quad[M_{r,\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{s}}]=\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}+\bar{s}}$, (2.9ah)
@@
$[M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=\left(\frac{1}{2}r-\frac{1}{2}s\right)Q^{(0,\frac{1}{2})}_{r+s,\bar{r}},\quad[M_{r,\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{s}}]=\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}+\bar{s}}$, (2.9ai)
@@
as well as Type II-I, which is the conjugate of Type I-II in (2.9ai). Owing to this conjugation relation, we will omit the discussion of Type II-I, as it follows analogously from the Type I-II case.
@@
The first case reduces to the extended super BMS4 algebras of Type I (2.9ac), Type II (2.9ad), or their combinations, via appropriate rotations of $Q^{(\frac{1}{2},0)}_{r}$ and $\tilde{Q}^{(\frac{1}{2},0)}_{s}$. In the second case, an additional bosonic generator arising from $\{\tilde{Q}^{(\frac{3}{2},0)},\tilde{Q}^{(\frac{3}{2},0)}\}$ would be required to satisfy the Jacobi identity of $\{Q^{(\frac{1}{2},0)}_{r},\tilde{Q}^{(\frac{3}{2},0)},\tilde{Q}^{(\frac{3}{2},0)}\}$. However, this generator lacks a well-defined physical motivation. Hence, no unextended super BMS4 algebra exists with two sets of $Q$s anticommuting to $L$. The same argument applies to the anti-holomorphic sector.
To conclude, there are six unextended magnetic super BMS4 algebras: chiral Type I and Type II, as well as nonchiral Type I-I, Type I-II, Type II-I, and Type II-II. In these algebras, the finite-dimensional subalgebra of Type I contains only spin-$\frac{1}{2}$ supercharges. However, it violates the ansatz that the anticommutators of supercharges yield translations. Therefore, none of the magnetic superalgebras are connected to the super Poincaré algebra.
@@ 3.1 Symmetry generators in terms of field modes
It is standard in QFT, particularly in 2D free CFTs and in BMS3 and BMS4 free theories, to express symmetry generators in terms of field modes [60, 61, 62, 63, 64, 65, 66]. The conventional approach to obtain the symmetry generators of a given free theory involves first computing the energy-momentum tensor in terms of the fields, then performing a Laurent or Fourier expansion of the fields according to their equations of motion (EOMs), and finally substituting this expansion to read off the generators. Here, since the symmetry algebra is already known, we adopt an inverse approach: we construct the symmetry generators using a specific ansatz for the field modes, and then build the corresponding field theory consistent with these mode expansions.
@@
$\varphi_{r},\phi_{s}\}=g^{l_{I}}_{l_{\varphi},l_{\phi}}(r,s)\delta_{r+s}I$ (2.9f)
@@
$L_{m},I_{n=0}]=(l_{I}*m-1*n)I_{m}=(0*m-1*0)I_{m}=0,$ (2.9g)
showing that additional generators $I_{m\neq 0}$ do not enter the theory as symmetry generators. Therefore, $\varphi$, $\phi$, and $I$ can all be viewed as charges in the theory, and equation (2.9f) implies that $g(r,s)\delta_{r+s}$ satisfies the constraint (2.8). The discrete solution relevant for this situation is
$f^{l_{I}}_{l_{\varphi},l_{\phi}}(r,s)=\delta_{r+s},\quad l_{\varphi}+l_{\phi}=-1,\quad l_{I}=0.$ (2.9h)
@@
$\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (2.9i)
@@
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}},$ (2.9j)
@@
In this construction, rule (2.9ea) is satisfied by the anti-holomorphic index of $L$, the holomorphic index of $\bar{L}$, and both indices of $M$, while rule (2.9eb) is satisfied by the holomorphic index of $L$ and the anti-holomorphic index of $\bar{L}$. In defining $L$ and $\bar{L}$, we consistently order the modes such that $\phi$ is always to the right of $\varphi$. This reproduces the BMS4 algebra (1.1) without introducing central terms. The ordering convention will be implicitly applied throughout the rest of this section. Furthermore, using the commutation relation (2.9i), one can verify that the transformation of the modes under the BMS4 algebra:
$[L_{m},\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{s,\bar{s}}]=\left(-\frac{3}{4}m-s\right)\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{m+s,\bar{s}}$, $[L_{m},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\left(-\frac{1}{4}m-s\right)\phi^{(-\frac{1}{4},-\frac{1}{4})}_{m+s,\bar{s}}$, (2.9k)
@@
$\xi^{(L)}_{n}=\left(-\frac{n+1}{2}tz^{n},-z^{(n+1)},0\right),$ (2.9l)
@@
$[L_{n},\phi]=-(\xi^{(L)}_{n})^{\mu}\partial_{\mu}\phi-2h\partial_{t}(\xi^{(L)}_{n})^{t}\phi+\frac{1}{2}\partial_{z}(\xi^{(L)}_{n})^{t}\phi_{z}+\frac{1}{2}\partial_{\bar{z}}(\xi^{(L)}_{n})^{t}\phi_{\bar{z}}+\cdots,$ (2.9m)
@@
Figure 1: Structure of Carrollian multiplet of $\phi$. The ellipses denote higher-order multiplet fields.
To meet equations (2.9m), the example discussed here can be identified as the free electric BMS4 scalar theory:
$\mathcal{L}=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi.$ (2.9n)
@@
$\phi(t,z,\bar{z})=\sum_{r,\bar{r}}\varphi_{r,\bar{r}}z^{-r-\frac{1}{4}}\bar{z}^{-\bar{r}-\frac{1}{4}}+t\phi_{r,\bar{r}}z^{-r-\frac{3}{4}}\bar{z}^{-\bar{r}-\frac{3}{4}}.$ (2.9o)
In this construction, the conformal dimensions of the field $\phi$ are $h=\bar{h}=\frac{1}{4}$, giving $\Delta=\frac{1}{2}$ and spin $j=0$. The commutation relation (2.9i) is consistent with the canonical commutation relation of the theory:
$\phi(t,z,\bar{z}),\partial_{0}\phi(t,z^{\prime},\bar{z}^{\prime})]=i\delta^{2}(z-z^{\prime}).$ (2.9p)
In the following subsections, we use the same method to construct supersymmetric BMS4 theories with the symmetries discussed in Section 2. For simplicity, we focus on scalar multiplets of different supersymmetries, with the scalar theory taken as (2.9n).

### v2

@@ 2.2 Electric super BMS4 algebra
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},1)}_{r,\bar{r}},Q^{(1,\frac{1}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},1)}_{r,\bar{r}},Q^{(1,\frac{1}{2})}_{s,\bar{s}}\}=\left(r-\frac{1}{2}s\right)\left(\frac{1}{2}\bar{r}-\bar{s}\right)M_{r+s,\bar{r}+\bar{s}},$ (2.23)
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(1,\frac{3}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(1,\frac{3}{2})}_{s,\bar{s}}\}=\left(r-\frac{1}{2}s\right)\frac{3}{2}\bar{r}M_{r+s,\bar{r}+\bar{s}},$ (2.24)
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{3}{2},0)}_{r,\bar{r}},Q^{(0,\frac{3}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{3}{2},0)}_{r,\bar{r}},Q^{(0,\frac{3}{2})}_{s,\bar{s}}\}=\frac{9}{4}s\bar{r}M_{r+s,\bar{r}+\bar{s}}.$ (2.25)
holomorphic (2.9b), anti-holomorphic (2.9c):
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(1,\frac{5}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{1}{2},0)}_{r,\bar{r}},Q^{(1,\frac{5}{2})}_{s,\bar{s}}\}=\left(r-\frac{1}{2}s\right)\bar{r}\left(\frac{5}{2}\bar{r}+\bar{s}\right)M_{r+s,\bar{r}+\bar{s}},$ (2.26)
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{3}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{3}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{s,\bar{s}}\}=\frac{3}{2}s\bar{r}\left(\frac{5}{2}\bar{r}+\bar{s}\right)M_{r+s,\bar{r}+\bar{s}}.$ (2.27)
holomorphic (2.9c), anti-holomorphic (2.9c):
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},Q^{(\frac{5}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{r,\bar{r}}\},\quad\{Q^{(\frac{5}{2},0)}_{r,\bar{r}},Q^{(0,\frac{5}{2})}_{s,\bar{s}}\}=s\left(r+\frac{5}{2}s\right)\bar{r}\left(\frac{5}{2}\bar{r}+\bar{s}\right)M_{r+s,\bar{r}+\bar{s}}.$ (2.28)
All of the above algebras contain finite-dimensional subalgebras. However, plugging in the coefficients, the finite-dimensional subalgebras are trivial with all $\{Q,Q\}=0$, except for (2.19), (2.20) and (2.23). Furthermore, (2.20) and (2.23) necessarily include spin-$\frac{3}{2}$ components. Such higher-spin fermionic generators are not compatible with the Haag–Łopuszański–Sohnius theorem[63], which allows only spin-$\frac{1}{2}$ supercharges in a consistent extension of the Poincaré algebra. Therefore, the only physically admissible super Poincaré algebra is the finite-dimensional subalgebra of (2.19).
@@ 2.3 Magnetic chiral super BMS4 algebras
$M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=\left(\frac{1}{2}r-\frac{1}{2}s\right)Q^{(0,\frac{1}{2})}_{r+s,\bar{r}},\quad\{Q^{(\frac{1}{2},0)}_{r},Q^{(\frac{1}{2},0)}_{s}\}=L_{r+s},\quad\{Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{s,\bar{s}}\}=M_{r+s,\bar{s}},$ (2.29)
@@
$M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=Q^{(1,\frac{1}{2})}_{r+s,\bar{r}},\quad\{Q^{(\frac{1}{2},0)}_{r},Q^{(\frac{1}{2},0)}_{s}\}=L_{r+s},\quad\{Q^{(\frac{1}{2},0)}_{r},Q^{(1,\frac{1}{2})}_{s,\bar{s}}\}=-\left(r-\frac{1}{2}s\right)M_{r+s,\bar{s}},$ (2.30)
In Type I, $M,Q^{(\frac{1}{2},0)}]\propto Q^{(0,\frac{1}{2})}$ satisfies (2.9b), whereas in Type II, $M,Q^{(\frac{1}{2},0)}]\propto Q^{(1,\frac{1}{2})}$ satisfies (2.9a). Unlike the electric case, the supercharges $Q^{(\frac{1}{2},0)}_{r}$ in (2.29) and (2.30) are not extended to $Q^{(\frac{1}{2},0)}_{r,\bar{r}}$, thereby avoiding the introduction of nonphysical generators $L_{m,\bar{m}}$. In contrast, the holomorphic index $r$ of $Q^{(0,\frac{1}{2})}_{r,\bar{r}}$ in (2.29) is required for the nonvanishing commutator $M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=Q^{(1,\frac{1}{2})}_{r+s,\bar{r}}$.
@@
$\{L_{n},\bar{L}_{\bar{n}},M_{r,\bar{r}},\tilde{M}_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{r,\bar{r}},Q^{(1,\frac{1}{2})}_{r,\bar{r}}\}$, (2.31)
@@
Observing (2.29), we note that $Q^{(\frac{1}{2},0)}$ and $Q^{(0,\frac{1}{2})}$ are not related by conjugation. This makes it clear that both (2.29) and (2.30) describe chiral superalgebras, as only the holomorphic Virasoro algebra $L_{n}$ appears in anticommutators of super charges. For this reason, (2.29) and (2.30) are considered as holomorphic in this paper. Anti-holomorphic counterparts $\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(\frac{1}{2},0)}_{r,\bar{r}}\}$ and $\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}}\}$ can be defined through conjugation, yielding super BMS4 algebra structures analogous to (2.29) and (2.30).
For a super BMS4 algebra containing both holomorphic and anti-holomorphic sectors, simply including the corresponding supercharges is insufficient. For instance, consider holomorphic supercharges $\{Q^{(\frac{1}{2},0)}_{r},Q^{(0,\frac{1}{2})}_{r,\bar{r}}\}$ and anti-holomorphic supercharges $\{\bar{Q}^{(0,\frac{1}{2})}_{\bar{r}},\bar{Q}^{(\frac{1}{2},0)}_{r,\bar{r}}\}$, both of Type I as in (2.29). The Jacobi identity for $\{Q^{(\frac{1}{2},0)},Q^{(\frac{1}{2},0)},\bar{Q}^{(\frac{1}{2},0)}\}$ is satisfied only if a bosonic generator $R^{(0,0)}$ is introduced, such that
$\{Q^{(\frac{1}{2},0)},\bar{Q}^{(\frac{1}{2},0)}\}\propto R^{(0,0)}.$ (2.32)
@@
$[M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=\left(\frac{1}{2}r-\frac{1}{2}s\right)Q^{(0,\frac{1}{2})}_{r+s,\bar{r}},\quad[M_{r,\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{s}}]=\left(\frac{1}{2}\bar{r}-\frac{1}{2}\bar{s}\right)\bar{Q}^{(\frac{1}{2},0)}_{r,\bar{r}+\bar{s}}$, (2.33)
@@
$[M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=Q^{(1,\frac{1}{2})}_{r+s,\bar{r}},\quad[M_{r,\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{s}}]=\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}+\bar{s}}$, (2.34)
@@
$[M_{r,\bar{r}},Q^{(\frac{1}{2},0)}_{s}]=\left(\frac{1}{2}r-\frac{1}{2}s\right)Q^{(0,\frac{1}{2})}_{r+s,\bar{r}},\quad[M_{r,\bar{r}},\bar{Q}^{(0,\frac{1}{2})}_{\bar{s}}]=\bar{Q}^{(\frac{1}{2},1)}_{r,\bar{r}+\bar{s}}$, (2.35)
@@
as well as Type II-I, which is the conjugate of Type I-II in (2.35). Owing to this conjugation relation, we will omit the discussion of Type II-I, as it follows analogously from the Type I-II case.
@@
The first case reduces to the extended super BMS4 algebras of Type I (2.29), Type II (2.30), or their combinations, via appropriate rotations of $Q^{(\frac{1}{2},0)}_{r}$ and $\tilde{Q}^{(\frac{1}{2},0)}_{s}$. In the second case, an additional bosonic generator arising from $\{\tilde{Q}^{(\frac{3}{2},0)},\tilde{Q}^{(\frac{3}{2},0)}\}$ would be required to satisfy the Jacobi identity of $\{Q^{(\frac{1}{2},0)}_{r},\tilde{Q}^{(\frac{3}{2},0)},\tilde{Q}^{(\frac{3}{2},0)}\}$. However, this generator lacks a well-defined physical motivation. Hence, no unextended super BMS4 algebra exists with two sets of $Q$s anticommuting to $L$. The same argument applies to the anti-holomorphic sector.
To conclude, there are six unextended magnetic super BMS4 algebras: chiral Type I and Type II, as well as nonchiral Type I-I, Type I-II, Type II-I, and Type II-II. All of these algebras admit finite-dimensional subalgebras. However, only in Type I-I and Type II-II we can define conjugation compatible with conjugation of Poincaré algebra, and the Gram matrices $\{Q,Q^{\dagger}\}$ for both cases are indefinite, which means the Hilbert spaces contain negative norm states. Therefore, none of the magnetic superalgebras are connected to the super Poincaré algebra.
@@ 3.1 Symmetry generators in terms of field modes
It is standard in QFT, particularly in 2D free CFTs and in BMS3 and BMS4 free theories, to express symmetry generators in terms of field modes [64, 65, 66, 67, 68, 69, 53, 70]. The conventional approach to obtain the symmetry generators of a given free theory involves first computing the energy-momentum tensor in terms of the fields, then performing a Laurent or Fourier expansion of the fields according to their equations of motion (EOMs), and finally substituting this expansion to read off the generators. Here, since the symmetry algebra is already known, we adopt an inverse approach: we construct the symmetry generators using a specific ansatz for the field modes, and then build the corresponding field theory consistent with these mode expansions.
@@
$\varphi_{r},\phi_{s}\}=g^{l_{I}}_{l_{\varphi},l_{\phi}}(r,s)\delta_{r+s}I$ (3.6)
@@
$L_{m},I_{n=0}]=(l_{I}*m-1*n)I_{m}=(0*m-1*0)I_{m}=0,$ (3.7)
showing that additional generators $I_{m\neq 0}$ do not enter the theory as symmetry generators. Therefore, $\varphi$, $\phi$, and $I$ can all be viewed as charges in the theory, and equation (3.6) implies that $g(r,s)\delta_{r+s}$ satisfies the constraint (2.8). The discrete solution relevant for this situation is
$f^{l_{I}}_{l_{\varphi},l_{\phi}}(r,s)=\delta_{r+s},\quad l_{\varphi}+l_{\phi}=-1,\quad l_{I}=0.$ (3.8)
@@
$\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{r,\bar{r}},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\frac{i}{4\pi^{2}}\delta_{r+s}\delta_{\bar{r}+\bar{s}},$ (3.9)
@@
$L_{m}=-4\pi^{2}i\sum_{s,\bar{s}}\left(s-\frac{1}{4}m\right)\varphi_{s,\bar{s}}\phi_{m-s,-\bar{s}},$ (3.10)
@@
In this construction, rule (3.5a) is satisfied by the anti-holomorphic index of $L$, the holomorphic index of $\bar{L}$, and both indices of $M$, while rule (3.5b) is satisfied by the holomorphic index of $L$ and the anti-holomorphic index of $\bar{L}$. In defining $L$ and $\bar{L}$, we consistently order the modes such that $\phi$ is always to the right of $\varphi$. This reproduces the BMS4 algebra (1.1) without introducing central terms. The ordering convention will be implicitly applied throughout the rest of this section. Furthermore, using the commutation relation (3.9), one can verify that the transformation of the modes under the BMS4 algebra:
$[L_{m},\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{s,\bar{s}}]=\left(-\frac{3}{4}m-s\right)\varphi^{(-\frac{3}{4},-\frac{3}{4})}_{m+s,\bar{s}}$, $[L_{m},\phi^{(-\frac{1}{4},-\frac{1}{4})}_{s,\bar{s}}]=\left(-\frac{1}{4}m-s\right)\phi^{(-\frac{1}{4},-\frac{1}{4})}_{m+s,\bar{s}}$, (3.11)
@@
$\xi^{(L)}_{n}=\left(-\frac{n+1}{2}tz^{n},-z^{(n+1)},0\right),$ (3.12)
@@
$[L_{n},\phi]=-(\xi^{(L)}_{n})^{\mu}\partial_{\mu}\phi-2h\partial_{t}(\xi^{(L)}_{n})^{t}\phi+\frac{1}{2}\partial_{z}(\xi^{(L)}_{n})^{t}\phi_{z}+\frac{1}{2}\partial_{\bar{z}}(\xi^{(L)}_{n})^{t}\phi_{\bar{z}}+\cdots,$ (3.13)
@@
Figure 1: Structure of Carrollian multiplet of $\phi$.
The ellipses denote higher-order multiplet fields.
To meet equations (3.13), the example discussed here can be identified as the free electric BMS4 scalar theory:
$\mathcal{L}=\frac{1}{2}\partial_{0}\phi\partial_{0}\phi.$ (3.14)
@@
$\phi(t,z,\bar{z})=\sum_{r,\bar{r}}\varphi_{r,\bar{r}}z^{-r-\frac{1}{4}}\bar{z}^{-\bar{r}-\frac{1}{4}}+t\phi_{r,\bar{r}}z^{-r-\frac{3}{4}}\bar{z}^{-\bar{r}-\frac{3}{4}}.$ (3.15)
In this construction, the conformal dimensions of the field $\phi$ are $h=\bar{h}=\frac{1}{4}$, giving $\Delta=\frac{1}{2}$ and spin $j=0$. The commutation relation (3.9) is consistent with the canonical commutation relation of the theory:
$\phi(t,z,\bar{z}),\partial_{0}\phi(t,z^{\prime},\bar{z}^{\prime})]=i\delta^{2}(z-z^{\prime}).$ (3.16)
In the following subsections, we use the same method to construct supersymmetric BMS4 theories with the symmetries discussed in Section 2. For simplicity, we focus on scalar multiplets of different supersymmetries, with the scalar theory taken as (3.14).